# Pathspace Kalman Filters with Dynamic Process Uncertainty for Analyzing Time-course Data

### v1

@@ 4 Efficient Bayesian Model Computations
For the KF, it is necessary that there be an internal model $m$ which generates a prediction at the next time-point $t$, $M_{t}\sim m(Z_{t-1})$ with defined mean and $\operatorname{\mathbb{E}}(M_{t})$ and variance $\operatorname{\sigma^{2}}(M_{t})$. Similarly, for the PKF to function, there must be an internal model $\operatorname{\boldsymbol{m}}$ which generates a predictions of the next filter iteration $i$ at each time-point $t$: $\operatorname{\boldsymbol{M}}_{t}^{i}=\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1},t)$. Notice that in the PKF the entire trajectory can be used to generate the model prediction. In order to expressly represent dynamics, we frame this as a Bayesian inference problem over the parameters of an ordinary differential equation. Consider the following ODE:
$\frac{\textrm{d}x}{\textrm{d}t}=g(x,k)t$ (10)
with solution over a time window $\Delta t=(t_{0},t_{f})$:
$x(k,\Delta t)=\int_{t_{0}}^{t_{f}}g(x;k)\textrm{d}t$ (11)
@@
We define $\operatorname{\boldsymbol{m}}(\zeta,t)=s(X,\Delta t,k)$ as a distribution that assigns solutions of the ODE (10) with parameter values $k$ a probability based on how well the solution fits all the data in a time window $\Delta t=(t_{0},t_{f})$ with $t_{0}\leq t\leq t_{f}$. Following common practice we let:
$s(\zeta,\Delta t,k)=\mathbb{P}(k\mid\zeta)\propto e^{-\mathcal{L}(x(k,\Delta_{t}),X_{\Delta t})}\pi(k)\$ (12)
@@
Here, $\pi(k)$ is a prior over the parameters, $\zeta_{\Delta t}$ denotes a vector of inputs in the range $(t_{0},t_{f})$ (these may come from the Filter $\operatorname{\boldsymbol{F}}$ or the data $Z$), and $\mathcal{L}$ is a loss function (such as the $L_{2}$ norm). In general, solving (12) for high dimensions of $k$ can be numerically challenging. We simplify this problem by choosing $\Delta t$ to be a window of 3 consecutive data points and using carefully chosen two parameter models $g$. This allows us to derive an analytic function $k_{2}=h(k_{1})$ constraining the parameters $k_{1}$ and $k_{2}$ of the model $g$ so it fits the first and third time-points exactly. We then varying a single parameter $k_{1}$ (which unique determines $k_{2}$ via $h$) to generate the distribution $s$ at the middle time-point $t$. This procedure is illustrated graphically in Figure (3). Intuitively $s$ can be thought of as a local distribution of splines consistent with the ODE (11). We describe this procedure more carefully in Appendix 8.3, including a modification necessary for start and end points of the trajectory. Different choices of $g$ are used for different kinds of data as described in the experiments section.
@@ 4.1 Interpretation of the Bayesian Method and Comparison to Process Uncertainty
In practice, the distribution of splines makes it easy to compute $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{M}}^{i}_{t})$ and $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})$ by sampling $s(\zeta,\Delta t,k)$ via MCMC in high dimensional cases or by scanning over an array of values $k$ to compute the relevant statistics directly in low dimensional cases. Here, the variance represents how tightly the ODE model can be fit to the input data which is typically the previous filter output $\operatorname{\boldsymbol{F}}^{i-1}$. By contrast, the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}^{i}$ is derived from the error between the model fit $\operatorname{\boldsymbol{M}}_{t}^{i}$ and the underlying data $Z_{t}$. These two sources of uncertainty are illustrated in Figure (3). The total variance at each point of time will be a combination of the variance of the underlying model ($\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})$), the process uncertainty ($\operatorname{\boldsymbol{Q}}_{t}^{i}$), and the variance in the data to which the model is fit ($\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i-1})$). Comparing the process uncertainty, the data variance, and the underlying model allows the PKF to detangle these different sources of uncertainties, which is infeasible using a traditional KF. We illustrate this analysis technique in the following section.
@@ 5 Experiments
In order to validate the PKF algorithm, we compared to a real KF using synthetic data because we needed access to the ground-truth state which is typically not present in real datasets. Then, to demonstrate the applicability of the PKF to real world problems, we used it to analyze a gene expression time-course dataset.
@@ 5.1 Kalman Filters on Synthetic Data
Data generation: Synthetic population dynamics data was used to compare standard KFs and the PKF models in order to have a ground truth answer to better evaluate the absolute performance of these models. Specifically, data were generated by simulating an ODE with time varying parameters and adding time-varying Gaussian noise to the output:
@@
Here $N(t)$ is the ground truth population with time varying birth and death rates, $k_{birth(t)}$ and $k_{death}(t)$. The data $Z(t)$ are $100$ samples per time-point drawn from a normal distribution $\mathcal{N}$ with mean $N(t)$ and time varying standard deviation $k_{noise}(t)$. $Z(t)$ are fed into the the different Kalman Filter models. Mean-square-error is computed by comparing the filter outputs to the ground truth data $N(t)$.
@@
Figure 4: Comparison of the KF and PKF on synthetic population dynamics data. A. The simulated dataset consisting of population dynamics generated from simulated a birth-death model and adding Gaussian noise. B. Parameters of the underlying model generating the simulated data. The growth and death rates change at the black vertical lines and the noise increases at the gray vertical line through the end of the time course, which are shown in all figures. C. Filter output a standard KF with $Q=1$. Notice that the filter estimate lags changes in the data generating process. D. Model and data weights for the same KF. E. Process uncertainty, filter variance, and mean-squared error for the same KF. F-G. Filter output and parameters for a standard KF with $Q=10$. Notice that increasing $Q$ decreases model weight. H. Process uncertainty, filter variance, and mean-squared error for the same KF. I. Filter output for the PKF shows no lag. J. Weights for the PKF show that after many iterations, the filter converges to a low variance estimate because the filter weight $\operatorname{\boldsymbol{u}}>\operatorname{\boldsymbol{w}},\operatorname{\boldsymbol{v}}$. K. Mean-square-error, filter variance, and process uncertainty for the PFK. Notice that the mean squared error is lowest for this model. Additionally, the process uncertainty spikes precisely when the data generating process changes.
Results for Synthetic Data: We compared the results of running a standard KF and the PKF on synthetic dataset described above. The internal model used by all the Kalman filter variants also takes the form of a birth-death process, namely the ODE:
@@
The parameters $k_{birth}$ and $k_{death}$ are fit dynamically over windows of 3 time-points from the underlying data as described in Appendix 8.4.1. To compare the different KF outputs, we use mean-squared-error as a metric:
@@
where $T$ are the total number of timepoints and either a KF or PKF can be used as the filter input into the expected value.
Columns 2 and 3 of Figure (4) illustrate how a standard KF with different process uncertainties fit this dataset. Although increasing the process uncertainty decreases the model weight globally, the traditional KF both still increase the model weight when data becomes noisier. We remark that the Kalman gain (data weight) is not monotonically decreasing because the data variance is a dynamic parameter in this example. Similarly, the filter variance is also non-monotonic in this example due to the use of a non-linear internal model.
The results of the PKF are illustrated in Column 4 of Figure (4). First, we note that the PKF has lower mean-squared-error than any of the other KFs at roughly 0.2% of the KF MSE. Second, we illustrate how the PKF is able to highlight temporal windows when the internal model’s dynamics are insufficient to explain the data. Notice that changes in the data generation parameters as indicated by the black dashed lines result in instantaneous changes in the process uncertainties demonstrating how the PKF can be used to highlight time windows when the internal model cannot fit the data. However, the PKF is able to quickly adapt to these changes and recalibrate its internal model at future time-points. Indeed, even after the noise is increased by a factor of 10 (dashed gray line), the filter is still able to fit the data well. Together, these observations lead to the following heuristic of comparing the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}$ and data variance $\operatorname{\sigma^{2}}({Z_{t}})$ as illustrated in Table (2).
@@
Low $\operatorname{\boldsymbol{Q}}_{t}$
High $\operatorname{\boldsymbol{Q}}_{t}$
Low $\operatorname{\sigma^{2}}(Z_{t})$
A) Accurate Model
@@
High $\operatorname{\sigma^{2}}(Z_{t})$
C) Accurate Model
@@
Convergence: In Figure (5) we show the numerical convergence of the algorithm on two different time-points. In Appendix (8.1.2) we prove that the PKF converges. This is seen by the rapid convergence of $\operatorname{\boldsymbol{Q}}_{t}^{i}$. The filter variance also converges towards 0, albeit very slowly. This behaviour may seem unexpected, however it occurs due to the feedback of the filter estimates into itself. In the limit of infinite iterations, the PKF has seen its own estimates increasingly more than it has seen the data or the model estimates which gives rise to decreasing filter variance. We emphasize that this behavior is not problematic as the most meaningful quantity in this framework is the process uncertainty and this converges to values representing the dynamic model uncertainty.
@@ 5.2 Application of the Pathspace Kalman Filter to Time-course Gene Expression Data
A prime motivation of this work is its application to biological time-courses, specifically to data obtained through high throughput acquisition methods such as bulk RNAseq and proteomics [23, 12]. In these methods, high dimensional molecular snapshots of biological processes are collected at individual times via destructive experiments. For example, in RNAseq experiments the expression of tens of thousands of genes may be measured simultaneously at each time-point but typically relatively few time-points are collected (usually at a resolution of several hours to days). The high dimensionality of this kind of data coupled with its temporal sparsity raises specific challenges. Firstly, biological systems are dynamic (not static) so one should expect them to change over time [10]. A crucial question is whether the observed changes themselves are expected or unexpected; dynamic process uncertainty allows for the detection of discrepancies between observations and an underlying prior model. Secondly, the scalability of any analysis method to function on high dimensional data is crucial because a central goal of analyzing this kind of data is ascribing novel function to poorly understood genes, for example as targets for therapeutic interventions [37]. In this section, we describe a model of gene expression suitable for detecting biologically important dynamics and apply it to a time-course gene expression dataset.
Dataset Description: In order to demonstrate the scalability of our approach on real world datasets, we used the PKF to analyze time-course gene expression measurements (RNAseq) of a circadian clock dataset in human cancer cell lines [44] with contains of roughly 1.8 million gene expression measurements consisting of 32337 gene expression measurements across two conditions and 14 time-points with two replicates per time-point and condition. Briefly, the mammalian circadian clock is an cell-autonomous oscillatory genetic circuit which regulates diverse cellular functions by synchronizing them to daylight patterns [31], [15]. For a brief description of the mammalian clock, refer to Appendix 8.4.2. In this dataset, the circadian clock is disrupted by activating the gene MyC [44] which causes the clock to be repressed. The effects of repressing the clock can be seen Figure (6A) which shows the dynamics of the core circadian oscillator gene BMAL1. The failure of BMAL1 to oscillate indicates the repression of the circadian clock. Figure (6B) additionally shows the filter output for BMAL1.
@@
Here, the expression rate ($k_{exp}$) and degradation rate ($k_{deg}$) of the transcript are assumed to be constant over short time windows–a hypothesis we call “constant regulation” which we discuss in more detail in Appendix 8.4.2. When the constant regulation model fails to accurately fit the data in a given window, it is indicative that the regulation of the gene is changing and is quantified by higher process uncertainty in the time window. This can be seen for BMAL1 in Figure (6C) where the process uncertainty is generally lower when the circadian clock is repressed.
@@
Additionally, we ran the PKF on all genes in this dataset and compared the two conditions. The average log ratio of process uncertainty to data variance for these genes is plotted against variance percentile in Figure (6 D). The circadian clock is known to regulate about 20% of the genome [31]. Our analysis picks up this signal by showing increased process uncertainty divided by data variance for the top $50^{\textrm{th}}$ percentile of most varying genes when the circadian clock is active compared to when it is repressed indicating that regulation is less dynamic when the clock is deactivated.

### v2

@@ 4 Efficient Bayesian Model Computations
A challenge of iteratively running the PKF in pathspace is the necessity of many evaluations of the internal model $\operatorname{\boldsymbol{m}}$. Notice that this model must generate a distribution of predictions at each time-point $\operatorname{\boldsymbol{M}}_{t}^{i}$ with a well defined mean and variance. Additionally, we desire that our internal model be representative of the underlying process. Therefore, we parameterize the model predictions by solutions to an ordinary differential equation with varying underlying parameters $k$ and its solution over a time window $\Delta t=(t_{0},t_{f})$:
$\frac{\textrm{d}x}{\textrm{d}t}=g(x,k)\quad\implies x(k,\Delta t)=\int_{t_{0}}^{t_{f}}g(x;k)\textrm{d}t$ (10)
We define the subtrajectory $\zeta_{\Delta t}$ of values in the range $\Delta t$ as the input to the model $\operatorname{\boldsymbol{m}}$. For $t\in\Delta t$, we define the PKF internal model as: $\operatorname{\boldsymbol{m}}(\zeta_{\Delta t},t)$ which is the distribution of the values of all paths $x(k,\Delta t)$ evaluated at time $t$. The probability $s(\zeta_{\Delta t},k)$ of each path as parameterized by $k$ is given by:
$s(\zeta_{\Delta t},k)=\mathbb{P}(k\mid\zeta_{\Delta t})\propto e^{-\mathcal{L}(x(k,\Delta t),\zeta_{\Delta t})}\pi(k)\$ (11)
@@
Here, $\pi(k)$ is a prior over the parameters and $\mathcal{L}$ is a loss function (such as the $L_{2}$ norm). In other words, $s(\zeta_{\Delta t},k)$ is a distribution that assigns solutions of the ODE (10) parameterized by $k$ a probability based on how well the solution fits all the data in the time window $\Delta t$.
There are three different weights $\operatorname{\boldsymbol{w}}$, $\operatorname{\boldsymbol{v}}$ and $\operatorname{\boldsymbol{u}}$ for the data, model, and filter, respectively. The symbols have been made bold to denote that they are vectors in both time $t$ (subscript) and filter iteration $i$ (superscript). For example $\operatorname{\boldsymbol{F}}^{i}=(\operatorname{\boldsymbol{F}}^{i}_{0},...,\operatorname{\boldsymbol{F}}^{i}_{T})$ with $T$ the number of time-points. Notice that the model samples $\operatorname{\boldsymbol{M}}^{i}_{t}\sim\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1},t)$ are now a function of the same time $t$ and the entire trajectory $\operatorname{\boldsymbol{F}}^{i-1}$ from the previous filter iteration instead of the single filter value $F_{t-1}$ at the previous time. Similarly, the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}^{i-1}$ is now a function of time and iteration instead of a constant. The prefactor $\operatorname{\boldsymbol{w}}^{i}_{t}+\operatorname{\boldsymbol{v}}^{i}_{t}$ in equation (8) is analogous to a Kalman Gain for the process uncertainty and weights its update proportional to the additional uncertainty the model requires in order to match the data.
@@
In general, solving (11) for high dimensions of $k$ can be numerically challenging but is possible using MCMC based approaches [foreman2013emcee]. We simplify this problem by carefully choosing two parameter models $g(x,k_{1},k_{2})$ which can be analytically solved. Then, we choose $\Delta t$ to be a window of 3 consecutive time points. This allows us to derive an analytic function $h$ such that $k_{2}=h(k_{1},\zeta_{\Delta t})$ which derives one parameter from the other given the data. We then vary a single parameter $k_{1}$ (which unique determines $k_{2}$ via $h$) to analytically compute the distribution $s$. This procedure is illustrated graphically in Figure (2). Intuitively $s$ can be thought of as a local distribution of splines consistent with the ODE (10). We describe this procedure more carefully in Appendix A.4. Different choices of $g$ are used for different kinds of data as described in the experiments section.
@@ 4.1 Interpretation of the Bayesian Method and Comparison to Process Uncertainty
In practice, the distribution of splines makes it easy to compute $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{M}}^{i}_{t})$ and $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})$ by sampling $s(\zeta_{\Delta t},k)$ via MCMC in high dimensional cases or by scanning over an array of values $k$ to compute the relevant statistics directly in low dimensional cases. Here, the variance represents how tightly the ODE model can be fit to the input data which is typically the previous filter output $\operatorname{\boldsymbol{F}}^{i-1}$. By contrast, the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}^{i}$ is derived from the error between the model fit $\operatorname{\boldsymbol{M}}_{t}^{i}$ and the underlying data $Z_{t}$. These two sources of uncertainty are illustrated in Figure (2). The total variance at each point of time will be a combination of the variance of the underlying model ($\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})$), the process uncertainty ($\operatorname{\boldsymbol{Q}}_{t}^{i}$), and the variance in the data to which the model is fit ($\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i-1})$). Comparing the process uncertainty, the data variance, and the underlying model allows the PKF to detangle these different sources of uncertainties, which is infeasible using a traditional KF. We illustrate this analysis technique in the following section and Table 2.
@@ 5 Experiments
In order to validate the PKF algorithm, we compare to several varieties of KF and Bayesian smoothing algorithms using synthetic data because we needed access to the ground-truth state which is typically not present in real datasets. Then, to demonstrate the applicability of the PKF to real world problems, we used it to analyze a gene expression time-course dataset.
@@ 5.1 Kalman Filters on Synthetic Data
Data generation: Synthetic population dynamics data were generated by simulating an ODE with time varying parameters and adding time-varying Gaussian noise to the output:
@@
Here $N(t)$ is the ground truth population with time varying birth and death rates, $k_{birth(t)}$ and $k_{death}(t)$. The data $Z(t)$ are $100$ samples per time-point drawn from a normal distribution $\mathcal{N}$ with mean $N(t)$ and time varying standard deviation $k_{noise}(t)$. $Z(t)$ are fed into the the different Kalman Filter models. Mean-square-error is computed by comparing the filter outputs to the ground truth data $N(t)$. This synthetic data is plotted in Figure 3A and B.
@@
Figure 3: Comparison of the adaptive non-linear univariate KF and PKF on synthetic population dynamics data. A. A synthetic dataset of population dynamics generated from simulating a birth-death model and adding Gaussian noise. B. Parameters of the underlying model generating the simulated data. The growth and death rates change at the black vertical lines and the noise increases at the gray vertical line through the end of the time course, as indicated in all panels. C. Filter output for an adaptive non-linear KF with $Q=1$. Notice that the filter estimate lags changes in the data generating process. D. Model and data weights for the same KF. E. Process uncertainty, filter variance, and mean-squared error for the same KF. F-G. Filter output and parameters for an adaptive non-linear KF with $Q=10$. Notice that increasing $Q$ decreases model weight. H. Process uncertainty, filter variance, and mean-squared error for the same KF. I. Filter output for the PKF shows no lag. J. Weights for the PKF show that after many iterations, the filter converges to a low variance estimate because the filter weight $\operatorname{\boldsymbol{u}}>\operatorname{\boldsymbol{w}},\operatorname{\boldsymbol{v}}$. K. Mean-squared error, filter variance, and process uncertainty for the PFK. Notice that the mean-squared error is lowest for this model. Additionally, the process uncertainty spikes precisely when the data generating process changes.
Results for Synthetic Data: We compared the results PKF with several KF varieties and Bayesian smoothers on synthetic dataset described above. These results are summarized in Table 3 and plotted in Figure 5. The internal model used by all the Kalman filter variants also takes the form of a birth-death process, namely the ODE:
@@
The parameters $k_{birth}$ and $k_{death}$ are fit dynamically over windows of 3 time-points from the underlying data as described in Appendix A.5.1. To compare the different KF outputs, we use mean-squared-error as a metric:
@@
where $T$ are the total number of time-points and either a KF or PKF can be used as the filter input into the expected value.
Columns 2 and 3 of Figure (3) illustrate how a KF with different process uncertainties fit this dataset. Although increasing the process uncertainty decreases the model weight globally, the KFs both still increase the model weight when data becomes noisier. We remark that the Kalman gain (data weight) is not monotonically decreasing because the data variance is a dynamic parameter in this example. Similarly, the filter variance is also non-monotonic in this example due to the use of a non-linear internal model.
The results of the PKF are illustrated in Column 4 of Table (3). First, we note that the PKF has lower mean-squared-error than any of the other KFs at roughly 0.2% of the KF MSE and roughly 10% of the best alternative Bayesian smoothing algorithm. Second, we illustrate how the PKF is able to highlight temporal windows when the internal model’s dynamics are insufficient to explain the data as illustrated in Figure (3). Notice that changes in the data generation parameters as indicated by the black dashed lines result in instantaneous changes in the process uncertainties demonstrating how the PKF can be used to highlight time windows when the internal model cannot fit the data. However, the PKF is able to quickly adapt to these changes and recalibrate its internal model at future time-points. Indeed, even after the noise is increased by a factor of 10 (dashed gray line), the filter is still able to fit the data well. Together, these observations lead to the following heuristic of comparing the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}$ and data variance $\operatorname{\sigma^{2}}({Z_{t}})$ as illustrated in Table 2.
@@
Low $\operatorname{\boldsymbol{Q}}_{t}$ High $\operatorname{\boldsymbol{Q}}_{t}$
Low $\operatorname{\sigma^{2}}(Z_{t})$ A) Accurate Model
@@
High $\operatorname{\sigma^{2}}(Z_{t})$ C) Accurate Model
@@ 5.2 Application of the Pathspace Kalman Filter to Time-course Gene Expression Data
A prime motivation of this work is its application to biological time-courses, specifically to data obtained through high throughput acquisition methods such as bulk RNAseq and proteomics [Wang2009, aebersold2003mass]. In these methods, high dimensional molecular snapshots of biological processes are collected at individual times via destructive experiments. For example, in RNAseq experiments the expression of tens of thousands of genes may be measured simultaneously at each time-point but typically relatively few time-points are collected (usually at a resolution of several hours to days). The high dimensionality of this kind of data coupled with its temporal sparsity raises specific challenges. Firstly, biological systems are dynamic (not static) so one should expect them to change over time [kitano2002systems]. A crucial question is whether the observed changes themselves are expected or unexpected; dynamic process uncertainty allows for the detection of discrepancies between observations and an underlying prior model. Secondly, the scalability of any analysis method to function on high dimensional data is crucial because a central goal of analyzing this kind of data is ascribing novel function to poorly understood genes, for example as targets for therapeutic interventions [paananen2020omics]. In this section, we describe a model of gene expression suitable for detecting biologically important dynamics and apply it to a time-course gene expression dataset.
Dataset Description: In order to demonstrate the scalability of our approach on real world datasets, we used the PKF to analyze time-course gene expression measurements (RNAseq) of a circadian clock dataset in human cancer cell lines [rebekah_clock] with contains of roughly 1.8 million gene expression measurements consisting of 32337 gene expression measurements across two conditions and 14 time-points with two replicates per time-point and condition. Briefly, the mammalian circadian clock is an cell-autonomous oscillatory genetic circuit which regulates diverse cellular functions by synchronizing them to daylight patterns [takahashi], [sgolden]. For a brief description of the mammalian clock, refer to Appendix A.5.2. In this dataset, the circadian clock is disrupted by activating the gene MyC [rebekah_clock] which causes the clock to be repressed. The effects of repressing the clock can be seen in Figure (4A) which shows the dynamics of the core circadian oscillator gene BMAL1. The failure of BMAL1 to oscillate indicates the repression of the circadian clock. Figure (4B) shows the filter output for BMAL1.
@@
Here, the expression rate ($k_{exp}$) and degradation rate ($k_{deg}$) of the transcript are assumed to be constant over short time windows–a hypothesis we call “constant regulation” which we discuss in more detail in Appendix A.5.2. When the constant regulation model fails to accurately fit the data in a given window, it is indicative that the regulation of the gene is changing and is quantified by higher process uncertainty in the time window. This can be seen for BMAL1 in Figure (4C) where the process uncertainty is generally lower when the circadian clock is repressed.
@@
Additionally, we ran the PKF on all genes in this dataset and compared the two conditions. The average log ratio of process uncertainty to data variance for these genes is plotted against variance percentile in Figure (4D). The circadian clock is known to regulate about 20% of the genome [takahashi]. Our analysis picks up this signal by showing increased process uncertainty divided by data variance for the top $50^{\textrm{th}}$ percentile of most varying genes when the circadian clock is active compared to when it is repressed indicating that regulation is less dynamic when the clock is deactivated.